# High-order interactions in quantum optomechanics: fluctuations, dynamics and thermodynamics

### v1

@@ Appendix B Modified perturbation theory
$\left(\hat{H}_{0}+\epsilon\hat{H}_{1}+\epsilon^{2}\hat{H}_{2}+\epsilon^{3}\hat{H}_{3}\right)\left(\lvert n^{(0)}\rangle+\epsilon\lvert n^{(1)}\rangle+\epsilon^{2}\lvert n^{(2)}\rangle+\epsilon^{3}\lvert n^{(3)}\rangle+...\right)$
$=\left(E_{n}^{(0)}+\epsilon E_{n}^{(1)}+\epsilon^{2}E_{n}^{(2)}+\epsilon^{3}E_{n}^{(3)}+...\right)\left(\lvert n^{(0)}\rangle+\epsilon\lvert n^{(1)}\rangle+\epsilon^{2}\lvert n^{(2)}\rangle+\epsilon^{3}\lvert n^{(3)}\rangle+...\right).$ (22)
@@ B.2 Second order in $\epsilon$
We now use the same technique utilized above to calculate the corrections at the second order in $\epsilon$ to both the eigenstates and the eigenenergies of the Hamiltonian. We therefore start from the second order Schrödinger equation
@@
where the last equivalence stems from the orthogonality between $\lvert n^{(1)}\rangle$ and $\lvert n^{(0)}\rangle$. The second order correction to the Hamiltonian eigenenergies is therefore given by
@@
Note that in the formula above we wrote $E_{n}^{(1)}$ explicitly via Eq. (24). The second order correction to the state is finally given by
@@
This formula differs from the standard one found in literature [63] because of the last contribution, which determines the transition between two eigenstates of the unperturbed Hamiltonian due to the second order interaction Hamiltonian $\hat{H}_{2}$.

### v2

@@ Appendix B Modified perturbation theory
[…]
@@ B.2 Second order in $\epsilon$
We now use the same technique utilized above to calculate the corrections at the second order in $\epsilon$ to both the eigenstates and the eigenenergies of the Hamiltonian. We therefore start from the second-order Schrödinger equation
@@
where the last equivalence stems from the orthogonality between $\lvert n^{(1)}\rangle$ and $\lvert n^{(0)}\rangle$. The second-order correction to the Hamiltonian eigenenergies is therefore given by
@@
Note that in the formula above we wrote $E_{n}^{(1)}$ explicitly via Eq. (33). The second-order correction to the state is finally given by
@@
This formula differs from the standard one found in literature [10] because of the last contribution, which determines the transition between two eigenstates of the unperturbed Hamiltonian due to the second-order interaction Hamiltonian $\hat{H}_{2}$.